Question:
(ON APPLYING BRAKES THE SPEED OF CAR REDUCES FROM 72KM/H TO 36KM/H IN 5 SECONDS WHAT IS THE RETARDATION OF CAR)
1
Expert's answer
2021-07-12T12:16:51-0400
"a=\frac{72-36}{3.6(5)}=2\frac{m}{s^2}"